The coefficient of linear expansion of steel is  

12 × 10-6/C°. A railroad track is made of individual rails of steel 1.0 km in length. By what length would these rails change between a cold day when the temperature is –10 °C and a hot day at 30 °C?


1
Expert's answer
2021-11-08T17:03:11-0500

"\Delta l=l_0\cdot\alpha\Delta t=1000\cdot12\cdot10^{-6}\cdot 40=0.48\ (m)"